How many molecules are there in 3 moles of aluminum chloride


1
Expert's answer
2021-05-14T05:56:42-0400

1 mole of Aluminum Chloride "= 6.022 \times 10^{23} \hspace{2mm}atoms"


3 moles of Aluminum Chloride "= 3 \times 6.022 \times 10^{23} = 18.06 \times 10^{23} \hspace{2mm}atoms"